# Discrete Guidance Matching: Exact Guidance for Discrete Flow Matching

### v1

@@ E.4 Classification of Discrete Guidance
$p_{s|t}(x_{s}|x_{t})=q_{s|t}(x_{s}|x_{t})$ for any $0\leq s<t\leq 1$ Continuous Rate-Based Guidance (Theorem 2 & [52]) $\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}{[u_{t}(z,x|{\mathbf{x}}_{1})]}\frac{p_{t}(x)q_{t}(z)}{p_{t}(z)q_{t}(x)}$
Discrete [76]& [64]& [37]
@@ E.5 Discussion on the Guidance Strength
In this subsection, we focus on the settings considered in [88, 86]. Our goal is to generate data from the guided distribution $p_{1}^{(\gamma)}(x|y)\propto p_{1}(x)p^{\gamma}(y|x)$, where $\gamma\geq 0$ is the guidance strength. Lemma 4.10 in [86] provides the comparison between contrastive energy prediction [48] and classifier guidance [16, 27]. Here, we compare the rate-based guidance with the predictor guidance [52] under different guidance strengths.
@@
allows us to sample from the guided distribution $p_{1}^{(\gamma)}$ with the probability path $p_{t}^{(\star,\gamma)}(x|y)\propto\sum_{x_{1}}p_{1}(x_{1})p_{1}^{\gamma}(y|x_{1})p_{t|1}(x|x_{1})$. The difference between the proposed rate-based guidance (equation 19) and the predictor guidance (equation 18) in discrete state space is similar to that between the exact guidance and the classifier guidance in continuous state space; see Lemma 4.10 of [86].
@@ G.1 Background and Motivation
In the diffusion language models [e.g. LLaDa, 51] and flow-based language models [e.g. Fudoki, 78], we can obtain the pretrained model. Recently, existing works have developed post-training methods to improve preference through DPO [60] or RLHF [55]. The main challenge is that the joint posterior predictor cannot be easily decoupled like AR models. To overcome this issue, some approximation strategies have been proposed for estimating log probability, such as empirical ELBO approximation [90] and mean-field approximation [87, 83]. However, these methods might introduce large approximation errors when the length of the sentence is large. Fortunately, our discrete guidance can provide a novel framework for alignment of diffusion and flow-based language models without approximation.
@@ G.2 Roadmap of Alignment for Flow-based Language Models
Designing conditional path and rate. Similar to [78], in practice, we take metric-induced probability path with kinetic-optimal conditional transition rate [65]:
@@
$~+\lambda\mathbb{E}_{t\sim\mathcal{U}([0,1]),{\mathbf{c}}\sim p_{\mathbf{c}},{\mathbf{o}}_{1}|{\mathbf{c}}\sim\pi^{*},{\mathbf{o}}_{t}|{\mathbf{o}}_{1}\sim q_{t|1}}{\Big[\sum_{d=1}^{\mathcal{D}}\log{\Big(\sum_{s\in\mathcal{S}}h^{d,\theta}_{t}(s,{\mathbf{o}}_{t},{\mathbf{c}})p_{1|t}^{d}(s|{\mathbf{c}},{\mathbf{o}}_{t})\Big)}-\log h^{d,\theta}_{t}({\mathbf{o}}_{1}^{d},{\mathbf{o}}_{t},{\mathbf{c}})\Big]}.$
@@ H.1 Implementation Details of Energy-Guided Sampling
For a fair comparison, we train the guidance models (conditional expectation) using the Bregman divergence (Equation 11 and Equation 20) without regularization. For the rate-based guidance and predictor guidance, we take the initial distribution $p_{0}=p_{0}^{(\gamma)}=\delta_{\mathbbm{m}}$ for efficient sampling, where $\mathbbm{m}$ is the mask state. For sampling with the posterior-based guidance, we consider both masked and uniform initial distributions. We used the mixture probability path and the associated transition rate (Equation 4) with the cosine time schedule $\kappa_{t}=\cos^{2}{[\frac{\pi}{2}(1-t)]}$ in our 2-D experiments, which is kinetic optimal as mentioned in [65].
H.2 Effectiveness of Regularization in a 2-D Experiment
@@ H.3 Additional 2-D Results
(b) Rate-Based (Masked, [52])
@@
Figure 7: Comparison of sampling results with different guidance schemes (64 sampling steps) in 2-D experiments. (a) The samples of the guided distribution $p_{1}^{(\gamma)}$ with different guidance strength $\gamma$; (b) the data sampled by the predictor guidance [52] with masked initial distribution; (c) the data sampled by the proposed rate-based guidance with masked initial distribution; (d) the data sampled by the proposed posterior-based guidance with masked initial distribution; (e) the data sampled by the proposed posterior-based guidance with uniform initial distribution.
@@
Rate-Based (Masked, [52])
@@
Rate-Based (Masked, [52])
@@
Rate-Based (Masked, [52])
@@
Rate-Based (Masked, [52])
@@
Rate-Based (Masked, [52])
@@
H.4 Ablation study of the regularization strength

### v2

@@ E.4 Classification of Discrete Guidance
$p_{s|t}(x_{s}|x_{t})=q_{s|t}(x_{s}|x_{t})$ for any $0\leq s<t\leq 1$ Continuous Rate-Based Guidance (Theorem 2 & Nisonoff et al. (2025)) $\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}{[u_{t}(z,x|{\mathbf{x}}_{1})]}\frac{p_{t}(x)q_{t}(z)}{p_{t}(z)q_{t}(x)}$
Discrete Vignac et al. (2023)& Schiff et al. (2025)& Li et al. (2024a)
@@ E.5 Discussion on the Guidance Strength
In this subsection, we focus on the settings considered in Zheng et al. (2023); Zhang et al. (2025). Our goal is to generate data from the guided distribution $p_{1}^{(\gamma)}(x|y)\propto p_{1}(x)p^{\gamma}(y|x)$, where $\gamma\geq 0$ is the guidance strength. Lemma 4.10 in Zhang et al. (2025) provides the comparison between contrastive energy prediction Lu et al. (2023) and classifier guidance Dhariwal and Nichol (2021); Ho and Salimans (2021). Here, we compare the rate-based guidance with the predictor guidance Nisonoff et al. (2025) under different guidance strengths.
@@
allows us to sample from the guided distribution $p_{1}^{(\gamma)}$ with the probability path $p_{t}^{(\star,\gamma)}(x|y)\propto\sum_{x_{1}}p_{1}(x_{1})p_{1}^{\gamma}(y|x_{1})p_{t|1}(x|x_{1})$. The difference between the proposed rate-based guidance (equation 19) and the predictor guidance (equation 18) in discrete state space is similar to that between the exact guidance and the classifier guidance in continuous state space; see Lemma 4.10 of Zhang et al. (2025).
@@ G.1 Background and Motivation
In the diffusion language models (e.g. LLaDa, Nie et al., 2025b) and flow-based language models (e.g. Fudoki, Wang et al., 2025), we can obtain the pretrained model. Recently, existing works have developed post-training methods to improve preference through DPO (Rafailov et al., 2023) or RLHF (Ouyang et al., 2022). The main challenge is that the joint posterior predictor cannot be easily decoupled like AR models. To overcome this issue, some approximation strategies have been proposed for estimating log probability, such as empirical ELBO approximation (Zhu et al., 2025) and mean-field approximation (Zhao et al., 2025; Yang et al., 2025). However, these methods might introduce large approximation errors when the length of the sentence is large. Fortunately, our discrete guidance can provide a novel framework for alignment of diffusion and flow-based language models without approximation.
@@ G.2 Roadmap of Alignment for Flow-based Language Models
Designing conditional path and rate. Similar to Wang et al. (2025), in practice, we take metric-induced probability path with kinetic-optimal conditional transition rate Shaul et al. (2025):
@@
$~+\lambda\mathbb{E}_{t\sim\mathcal{U}([0,1]),{\mathbf{c}}\sim p_{\mathbf{c}},{\mathbf{o}}_{1}|{\mathbf{c}}\sim\pi^{*},{\mathbf{o}}_{t}|{\mathbf{o}}_{1}\sim q_{t|1}}{\Big[\sum_{d=1}^{\mathcal{D}}\log{\Big(\sum_{s\in\mathcal{S}}h^{d,\theta}_{t}(s,{\mathbf{o}}_{t},{\mathbf{c}})p_{1|t}^{d}(s|{\mathbf{c}},{\mathbf{o}}_{t})\Big)}-\log h^{d,\theta}_{t}({\mathbf{o}}_{1}^{d},{\mathbf{o}}_{t},{\mathbf{c}})\Big]}.$ (22)
@@ H.1 Implementation Details of Energy-Guided Sampling
For a fair comparison, we train the guidance models (conditional expectation) using the Bregman divergence (Equation 11 and Equation 20) without regularization. For the rate-based guidance and predictor guidance, we take the initial distribution $p_{0}=p_{0}^{(\gamma)}=\delta_{\mathbbm{m}}$ for efficient sampling, where $\mathbbm{m}$ is the mask state. For sampling with the posterior-based guidance, we consider both masked and uniform initial distributions. We used the mixture probability path and the associated transition rate (Equation 4) with the cosine time schedule $\kappa_{t}=\cos^{2}{[\frac{\pi}{2}(1-t)]}$ in our 2-D experiments, which is kinetic optimal as mentioned in Shaul et al. (2025).
Our pretrained model and guidance models are SiLU networks with 3 hidden layers with dimension 256. For training, we use 100,000 datapoints from source data distribution. The optimizer is Adam with learning rate 1e-4.
H.2 Implementation Details of Multimodal Tasks
To match the conditional probability path used in Wang et al. (2025), we use metric-induced path with metric $d(f(x),f(z))=\|{\tilde{x}-\tilde{z}}\|^{4}$, where $\tilde{x},\tilde{z}$ are normalized token embeddings which are taken from the original text embedding layer of Janus-Pro-7B and the image embeddings of LlamaGen; for time schedule $\beta_{t}$, we set $\beta_{t}=3(\frac{t}{1-t})^{0.9}$ as suggested in Shaul et al. (2025).
Our guidance model is a network composed of 6 LLaDA blocks (Nie et al., 2025c) with hidden dimension 768, a time embedding mapping $\mathbb{R}$ to a hidden dimension of 768, a token embedding mapping each token into a 768-dimensional space.
H.3 Effectiveness of Regularization in a 2-D Experiment
@@ H.4 Additional 2-D Results
(b) Rate-Based (Masked, Nisonoff et al. (2025))
@@
Figure 7: Comparison of sampling results with different guidance schemes (64 sampling steps) in 2-D experiments. (a) The samples of the guided distribution $p_{1}^{(\gamma)}$ with different guidance strength $\gamma$; (b) the data sampled by the predictor guidance (Nisonoff et al., 2025) with masked initial distribution; (c) the data sampled by the proposed rate-based guidance with masked initial distribution; (d) the data sampled by the proposed posterior-based guidance with masked initial distribution; (e) the data sampled by the proposed posterior-based guidance with uniform initial distribution.
@@
Rate-Based (Masked, Nisonoff et al. (2025))
@@
Rate-Based (Masked, Nisonoff et al. (2025))
@@
Rate-Based (Masked, Nisonoff et al. (2025))
@@
Rate-Based (Masked, Nisonoff et al. (2025))
@@
Rate-Based (Masked, Nisonoff et al. (2025))
@@
H.5 Ablation study of the regularization strength
@@
LLM Usage
In our multimodal understanding experiments, we follow the standard practice of using LLM-as-a-judge to assess rewards and evaluations based on the question, ground-truth answer, and model response. LLMs are further used solely for polishing the writing. Importantly, no novel ideas, analyses, or discoveries are contributed by LLMs.